OpenStudy (britbrat1997):

The sum of the ages of Cesar and April is 79 years. 11 years ago, Cesar's age was 2 times April's age. How old is Cesar now?

OpenStudy (boldjon):

CESAR IS DEAD

OpenStudy (boldjon):

THE GREAT JULIUS CESAR ID DEAD

OpenStudy (isaidavila):

Is just a system of equations C+A=79 C-11=2(A-11) You can solve that!!!